A buffer is prepared by adding 1.00 L of 1.0 M HCl to 750 mL of 1.5 M NaHCOO.

What is the pH of this buffer? Ka

= 1.7
Answer #1
Concepts and reason

The concept used to solve this problem is based on buffer solutions.

The pH of an acidic buffer solution is determined with the help of Henderson-Hasselbalch equation.

Fundamentals

Buffer solutions are those which resists change in its pH when a small amount of an acid or a base are added to it. The Henderson-Hasselbalch equation is shown as follows:

pH = pK+log
salt
(acid]

Here, is concentration of salt and concentration of weak acid.

The dissociation of the acid is shown as follows:

initially
At equilibrium
HCOO + HCI = Cl + HCOOH
1.125 1 0 0
(1.125-1) (1-1) 1 1

Now, convert volume into liters as follows:

Y-750 ml IL
V = 750 mL
(1000 ml)
= 0.750 L

Now, total volume is calculated as follows:

V =V,+V2

Substitute 0.750 L
for and 1.00 L for .

V = 0.750 L +1.00 L
= 1.75 L

Now calculate the concentration in terms of molarity as follows:

[HCOOH) = 1,75L
u_ 1 mol
= 0.57 M

And,

(1.125-1) mol
1.75 L
0.125 mol
1.75 L
= 0.071 M

Now, Henderson-Hasselbalch equation is as follows:

pH = pK+log
salt
(acid]

Substitute 3.77 for , 0.125 for and 1 for .

0.125
pH = 3.77 +log
= 3.77 +log 0.125
= 3.77+(-0.90)
= 2.87

Ans:

The pH of the given buffer is 2.87.